# LLMs Are Biased Towards Output Formats! Systematically Evaluating and Mitigating Output Format Bias of LLMs

### v1

@@ C.2 Wrapping Discussions
Gemma-7B-it (Systematic-$F1$) 6.62 / 11.74 4.68 / 0.64 4.75 / 9.79 1.77 / 1.21 0.00 / 0.00 2.58 / 4.08 0.0 / 0.0
Gemma-7B-it (FI) 38.00 / 39.50 52.50 / 66.00 58.50 / 67.50 8.00 / 6.50 0.00 / 0.00 12.50 / 18.50 0.0 / 0.0
@@
Table 6: Wrapping output format results. Red results are unreliable results measured by Theorem 3.1 i.e., inequality Eq. 6 does not hold.
@@
Gemma 31.09 20.11 24.77 23.39 14.61 33.86 11.63 56.33
Mistral 34.06 42.43 27.91 37.44 32.83 35.49 30.90 18.83
ChatGPT 32.47 33.40 32.95 38.16 38.49 38.40 28.69 12.26
Average 32.54 31.98 28.54 33.00 28.64 35.92 23.74 13.55
@@
MMLU 49.42 56.72 57.85 54.91 52.24 48.77 63.63 23.26
BBH 44.77 36.85 26.50 40.00 26.40 30.87 18.67 70.16
GSM8K 17.00 15.35 15.35 24.56 21.10 24.16 19.00 12.97
HotpotQA 17.62 30.46 24.46 23.91 15.46 32.21 15.19 41.40
FairytaleQA 28.40 27.10 30.18 26.12 22.29 31.70 18.84 17.28
Average 31.44 33.29 30.87 33.90 27.50 33.57 27.06 6.88
Table 7: Avg. estimated true accuracy results of benchmarks across models (Gemma-7B-it, Mistral, ChatGPT) and datasets (MMLU, BBH) with seven Wrapping format (Special Character, Bolding, Italicizing, Brackets, Parentheses, Placeholder, Quoting). This table is the distillation result of averaging all the estimated true performance scores across models and benchmarks in Tab. 6.
@@
The FI scores over formats are provided in Tab. 8. Overall, LLMs exhibit significant format-following bias across formats with a variance of FI scores of $297.28\%^{2}$. Among the models, ChatGPT follows the instructions best with average FI Score $85.01\%$. The “Special Character” wrapping format has the highest FI score of $73.34\%$. Following it is the “Placeholder” wrapping format also shows a high FI score of $68.37\%$, suggesting it is another effective format for ensuring instruction adherence. In contrast, the “Quoting” wrapping format has the lowest FI score of $17.06\%$. This significant drop compared to other formats suggests that quoting is the least effective method for wrapping instructions, possibly causing confusion or misinterpretation by the models. Below we present two examples of Gemma and Mistral failing to follow the format instructions:
@@
The average of estimated True Accuracy scores across models and tasks for the ZS and ZS-CoT prompting are shown in Tab. 9. For the majority of the wrapping methods (“Special Character”, “Bolding”, “Italicizing”, and “Brackets”), the ZS-CoT model generally shows higher or comparable performance to the ZS model. The “Brackets” shows a significant improvement when using ZS-CoT, with a jump from 31.05% (ZS) to 36.00% (ZS-CoT). However, the $BiasF_{o}$ metric shows a considerable divergence between $4.89\%^{2}$ (ZS) and $11.65\%^{2}$ (ZS-CoT). This depicts that while applying CoT may improve accuracy in certain methods, it does not generally reduce format bias.
@@
Zero-shot 30.57 32.76 30.30 31.05 27.38 34.42 28.50 4.89
Zero-shot Chain-of-Thought 27.99 34.32 32.33 36.00 27.05 31.02 26.58 11.65
@@
Overall, $80\%$ of the $EstTrue$ results ($169/210$) are reliable. Gemma-7B-it shows mixed reliability, with some red-marked scores indicating unreliable results, particularly in the “Quoting” format. This is because Gemma failed to follow the quoting instruction to quote the final answer. Mistral-7B-it exhibits similar variability, with some unreliable scores in “Quoting” and “Placeholder” formats. ChatGPT generally demonstrates mostly reliable results, with only 1 quoting result unreliable.
@@ C.3 List Discussions
Llama-3.1-8B-it (EstTrue-$F1$) - / - - / - - / - - / -
@@
Llama-3.1-8B-it (EstTrue-$F1$) - / - - / - - / - - / -
@@
Table 10: List output format results. Red results are unreliable results measured by Theorem 3.1 i.e., inequality Eq. 6 does not hold.
@@
SemEval2017 21.58 8.09 10.43 16.29 27.58
SciDocsRR 41.66 30.43 28.91 48.78 67.08
Average 31.62 19.26 19.67 32.54 39.91
@@
From Tab. 12, the “Newline list” format has the highest FI score of $66.05\%$, while the FI scores of other formats are around $48.6\%$. Among the models, ChatGPT follows the instructions best with FI score $83.27\%$. Specifically, ChatGPT excels at following instructions for all formats. In comparison, Mistral performs poorly with “Special character list”, while Gemma struggles the most with “Python” and “Bullet point list”. Explaining these phenomena, we observe that Gemma often misinterprets the task of converting textual samples into lists of numbers as a coding task, resulting in answers that do not match the desired format or content. We take an example in dataset SciDocsRR Section D.1 with detailed prompt as in Section E.3
@@
The results, detailed in Tab. 13 indicate that prompting with ZS-CoT substantially enhances model performance across various formats. Moreover, ZS-CoT effectively reduces format bias, as evidenced by the $BiasF_{o}$ metric decreasing from $46.88\%^{2}$ to $33.69\%^{2}$. From this, we conclude that CoT reduces format bias.
@@
Zero-shot 27.76 13.67 14.73 27.98 46.88
Zero-shot Chain-of-Thought 35.47 24.85 24.62 37.09 33.69
@@
From Tab. 10, $73\%(35/48)$ of the $EstTrue$ results are reliable. However, some scores of Gemma-7B-it ($8/16$) and Mistral-7B-it ($5/16$) on these benchmarks are red-marked, indicating unreliable results of this model. In contrast, the ChatGPT’s results are perfectly reliable.
@@ C.4 Mapping Discussions
Table 14: Mapping output format results. Red results are unreliable results measured by Theorem 3.1 i.e., inequality Eq. 6 does not hold.
@@
Easy 25.92 11.55 18.74 51.62
Medium 21.69 16.82 19.26 5.92
Hard 26.53 9.15 17.84 75.51
@@
From this Tab. 16, noticeably, there is a significant difference in the FI scores of two mapping formats (“JSON” and “YAML”). The average FI scores of the “JSON” mapping format is $66.07\%$, which is exceptionally higher (double time higher) than the one of the “YAML” mapping format ($34.62\%$), resulting in high format bias ($BiasF_{o}$ = $494.64\%^{2}$). Overall, among models, ChatGPT yields the best result of $90.40\%$. Mistral fails to follow “YAML” format instruction, e.g. for the question in dataset SciREX Easy Section D.1 with detailed promt as in Section E.4
@@
Gemma 20.71 12.62 16.66 32.68
Mistral 82.82 5.13 43.98 3018.00
ChatGPT 94.69 86.11 90.40 36.82
Avg. 66.07 34.62 50.35 494.64
@@
Zero-shot 25.59 13.70 35.30
Zero-shot Chain-of-Thought 23.84 11.31 39.29
@@
From Tab. 14, $75\%$ of the $EstTrue$ results are reliable. The reliability of the results in the mapping output format shows variability across different models and formats. Noticeably, “YAML” mapping format results are less reliable than “JSON” ones. On the other hand, ChatGPT illustrates its high reliability in all mapping formats while Mistral-7B-it and Gemma-7B-it are opposite, and all the results in the “YAML” mapping format of these models are unreliable.
@@ C.5 Mitigating Format Bias Results
Table 18: Supplementing demonstrations, repeating format instructions, and extra fine-tuning with formats’ data reduce format bias. Performance of ChatGPT on MMLU. All results are reliably measured by Theorem 3.1 i.e., inequality Eq. 6 holds.
In this section, we present the numerical results of our proposed techniques for mitigating format biases using ChatGPT on MMLU, as shown in Tab. 18.

### v2

@@ C.2 Wrapping Discussions
Gemma-7B-it (Systematic-$F1$) 6.62 / 11.74 4.68 / 0.64 4.75 / 9.79 1.77 / 1.21 0.00 / 0.00 2.58 / 4.08 0.00 / 0.00
Gemma-7B-it (FI) 38.00 / 39.50 52.50 / 66.00 58.50 / 67.50 8.00 / 6.50 0.00 / 0.00 12.50 / 18.50 0.00 / 0.00
@@
Table 6: Wrapping results. Red results are unreliable results measured by Theorem 3.1 i.e., inequality Eq. 6 does not hold.
@@
Gemma 17.69 22.23 28.24 23.35 9.86 25.88 22.25 31.31
Mistral 32.87 29.33 31.05 28.46 26.02 33.23 24.64 9.20
LLama 44.41 30.16 28.98 36.78 35.67 45.69 18.85 74.86
ChatGPT 40.15 44.91 32.65 45.24 42.72 43.82 32.57 26.03
Average 33.78 31.65 30.23 33.46 28.57 37.15 24.58 14.14
@@
MMLU 50.56 52.48 60.59 55.18 48.09 53.24 58.65 16.58
BBH 47.13 34.96 24.68 38.98 29.77 38.59 26.32 54.58
GSM8K 25.12 16.40 14.17 25.55 25.47 26.35 14.25 28.51
HotpotQA 18.50 27.35 19.75 22.59 16.27 29.41 11.39 33.66
FairytaleQA 32.10 31.36 32.78 30.79 28.11 35.40 14.13 42.16
Average 34.68 32.51 30.39 34.62 29.54 36.60 24.95 13.39
Table 7: Avg. estimated true accuracy results of benchmarks across models and datasets with seven Wrapping formats (Special Character, Bolding, Italicizing, Brackets, Parentheses, Placeholder, Quoting). This table is the distillation result of averaging all the estimated true performance scores across models and benchmarks in Tab. 6.
@@
The FI scores over formats are provided in Tab. 8. Overall, LLMs exhibit significant format-following bias across formats with a variance of FI scores of $345.85\%^{2}$. Among the models, ChatGPT follows the instructions best with an average FI Score of $85.01\%$. The “Special Character” wrapping format has the highest FI score of $75.05\%$. Following it is the “Placeholder” wrapping format also shows a high FI score of $72.47\%$, suggesting it is another effective format for ensuring instruction adherence. In contrast, the “Quoting” wrapping format has the lowest FI score of $18.55\%$. This significant drop compared to other formats suggests that quoting is the least effective method for wrapping instructions, possibly causing confusion or misinterpretation by the models. Below we present two examples of Gemma and Mistral failing to follow the format instructions:
@@
The average of estimated True Accuracy scores across models and tasks for the ZS and ZS-CoT prompting are shown in Tab. 9. For the majority of the wrapping methods (“Special Character”, “Bolding”, “Italicizing”, and “Brackets”), the ZS-CoT model generally shows higher or comparable performance to the ZS model. The “Italicizing” shows a significant improvement when using ZS-CoT, with a jump from 28.28% (ZS) to 32.19% (ZS-CoT). However, the $BiasF_{o}$ metric shows a considerable difference between $11.39\%^{2}$ (ZS) and $16.22\%^{2}$ (ZS-CoT). This depicts that while applying CoT may improve accuracy in certain methods, it does not generally reduce format bias.
@@
Zero-shot 34.31 30.66 28.28 33.82 28.77 35.93 26.09 11.39
Zero-shot Chain-of-Thought 33.25 32.65 32.19 33.09 28.36 38.37 24.46 16.22
@@
Overall, $80\%$ of the $EstTrue$ results are reliable. Gemma-7B-it shows mixed reliability, with some red-marked scores indicating unreliable results, particularly in the “Quoting” format. This is because Gemma failed to follow the quoting instruction to quote the final answer. Mistral-7B-it exhibits similar variability, with some unreliable scores in “Quoting” and “Placeholder” formats. ChatGPT generally demonstrates mostly reliable results, with only 1 quoting result unreliable.
@@ C.3 List Discussions
Llama-3.1-8B-it (EstTrue-$F1$) 32.10 / 0.00 38.80 / 0.00 0.00 / 0.00 35.78 / 0.00
Llama-3.1-8B-it (Systematic-$F1$) 6.26 / 0.00 3.88 / 0.00 0.00 / 0.00 35.78 / 0.00
Llama-3.1-8B-it (FI) 19.50 / 0.00 10.00 / 0.00 0.00 / 0.00 100.00 / 0.00
@@
Llama-3.1-8B-it (EstTrue-$F1$) 34.82 / 0.00 0.22 / 0.00 0.00 / 0.00 0.00 / 10.17
Llama-3.1-8B-it (Systematic-$F1$) 33.95 / 0.00 0.22 / 0.00 0.00 / 0.00 0.00 / 10.17
Llama-3.1-8B-it (FI) 97.50 / 0.00 100.00 / 100.00 0.00 / 75.50 100.00 / 100.00
@@
Table 10: List results. Red results are unreliable results measured by Theorem 3.1 i.e., inequality Eq. 6 does not hold.
@@
SemEval2017 20.54 6.09 7.82 13.49 31.86
SciDocsRR 35.25 27.67 21.68 41.06 54.12
Average 27.90 16.88 14.75 27.27 46.97
@@
From Tab. 12, the “Newline list” format has the highest FI score of $68.29\%$, while the FI scores of other formats are around $40-50\%$. Among the models, ChatGPT follows the instructions best with FI score $83.27\%$. Specifically, ChatGPT excels at following instructions for all formats. In comparison, Mistral performs poorly with “Special character list”, while Gemma struggles the most with “Python” and “Bullet point list”. Explaining these phenomena, we observe that Gemma often misinterprets the task of converting textual samples into lists of numbers as a coding task, resulting in answers that do not match the desired format or content. We take an example in dataset SciDocsRR Section D.1 with detailed prompt as in Section E.3
@@
The results, detailed in Tab. 13 indicate that prompting with ZS-CoT substantially enhances model performance across various formats. Moreover, ZS-CoT effectively reduces format bias, as evidenced by the $BiasF_{o}$ metric decreasing from about $54\%^{2}$ to $22\%^{2}$. From this, we conclude that CoT reduces format bias.
@@
Zero-shot 29.19 15.13 11.05 25.45 54.48
Zero-shot Chain-of-Thought 26.61 18.64 18.46 29.09 22.40
@@
From Tab. 10, $67\%$ of the $EstTrue$ results are reliable. However, some scores of Gemma-7B-it and LLama on these benchmarks are red-marked, indicating unreliable results of this model. In contrast, the ChatGPT’s results are perfectly reliable.
@@ C.4 Mapping Discussions
Table 14: Mapping results. Red results are unreliable results measured by Theorem 3.1 i.e., inequality Eq. 6 does not hold.
@@
Easy 25.53 14.61 20.07 29.79
Medium 25.74 19.96 22.85 8.36
Hard 29.77 13.54 21.66 65.87
@@
From this Tab. 16, noticeably, there is a significant difference in the FI scores of two mapping formats (“JSON” and “YAML”). The average FI scores of the “JSON” mapping format is $60.98\%$, which is much higher than “YAML” mapping format ($43.96\%$). Overall, among models, ChatGPT yields the best result of $90.40\%$. Mistral fails to follow “YAML” format instruction, e.g. for the question in dataset SciREX Easy Section D.1 with detailed promt as in Section E.4
@@
Gemma 20.71 12.62 16.66 16.34
Mistral 82.82 5.13 43.98 1509.00
Llama 45.70 71.97 58.84 172.46
ChatGPT 94.69 86.11 90.40 18.41
Avg. 60.98 43.96 52.47 72.45
@@
Zero-shot 28.02 16.79 31.55
Zero-shot Chain-of-Thought 26.01 15.29 28.73
@@
From Tab. 14, $77\%$ of the $EstTrue$ results are reliable. The reliability of the results in the mapping output format shows variability across different models and formats. Noticeably, “YAML” mapping format results are less reliable than “JSON” ones. On the other hand, ChatGPT illustrates its high reliability in all mapping formats while Mistral-7B-it and Gemma-7B-it are opposite, and all the results in the “YAML” mapping format of these models are unreliable.
@@ C.5 Mitigating Format Bias Results
No demo (Zero-shot)
16 Gemma-2B (EstTrue-Acc) 30.64 6.88 1.90 0.0 0.0 8.41 0.0 104.81
17 Gemma-2B (Systematic) 23.26 4.32 0.78 0.0 0.0 8.32 0.0 63.00
18 Gemma-2B (FI) 76.24 62.83 41.06 0.0 0.0 98.91 0.0 39.86 1443.70
Repeat format prompt quintice
19 Gemma-2B (EstTrue-Acc) 58.40 0.61 32.26 0.00 0.00 55.52 0.00 636.50
20 Gemma-2B (Systematic) 56.00 0.40 14.00 0.00 0.00 54.00 0.00 576.77
21 Gemma-2B (FI) 95.89 65.35 43.40 0.00 0.00 97.27 0.00 43.13 1684.16
1 demo
22 Gemma-2B (EstTrue-Acc) 35.57 34.45 34.84 31.76 0.00 26.35 34.34 140.95
23 Gemma-2B (Systematic) 32.03 34.45 25.25 26.81 0.00 6.07 27.49 84.32
24 Gemma-2B (FI) 90.03 99.98 72.47 84.41 0.00 23.03 80.05 74.99 611.87
5 demos
25 Gemma-2B (EstTrue-Acc) 38.56 38.90 38.49 36.72 36.52 38.97 37.25 0.94
26 Gemma-2B (Systematic) 38.56 38.90 38.49 36.72 36.08 38.29 37.20 0.99
27 Gemma-2B (FI) 100.00 100.00 100.00 100.00 98.78 98.26 99.86 99.57 0.45
Table 18: Supplementing demonstrations, repeating format instructions, and extra fine-tuning with formats’ data reduce format bias. Performance of ChatGPT and Gemma-2B-it (without CoT) on MMLU. All results are reliably measured by Theorem 3.1.
In this section, we present the numerical results of our proposed techniques for mitigating format biases using ChatGPT on MMLU, as shown in Tab. 18 indexes 1-15.
@@
For Gemma-2B-it model indexed 16-27, we observe that:
  1. 1.
    Repeating format instruction (Indexes 19-21). This approach does not resolve Gemma-2B-it’s inability to adhere to the “Brackets” and “Parentheses” instructions. However, disregarding these two formats, the strategy effectively reduces format bias in the model’s performance.
  2. 2.
    Demonstrations (Indexes 22-27). Demonstrations significantly mitigate format bias, as evidenced by higher FI scores and reduced variance across format performances.